OpenStudy (anonymous):

1. Choose an integer 2. Multiply it by 3 3. Add 6 to the product. 4. Now divide by 3. 5. Subtract 2. 6. The result will be the integer you started with. Why? How can you explain why the result is always the integer you started with?

OpenStudy (anonymous):

let that integer be n all the steps: (((n * 3) + 6) / 3) - 2 no simplify: [(3n + 6) / 3] - 2 = [n + 2] - 2 = n

OpenStudy (anonymous):

I have followed the steps already.

OpenStudy (anonymous):

so what is the problem ? we got n back.. when you simplify it we see that every thing disappears but the integer

OpenStudy (anonymous):

How can you explain why the result is always the integer you started with?

OpenStudy (anonymous):

as you can see all the operations we did are cancelled eventually ..

OpenStudy (anonymous):

Just a moment please.

OpenStudy (anonymous):

...I kind of get it.

OpenStudy (anonymous):

Let's say that the integer is 15. 15 times 3 = 45. 45 + 6 = 51 51 divided by 3 = 17 17 - 2 = 15

OpenStudy (anonymous):

I don't quite understand how you simplified it.

OpenStudy (anonymous):

the first step is 15 * 3 (dont write 45 leave it as 15 * 3 ) the second is 15 * 3 + 6 third is (15 * 3 + 6) / 3 forth (15 * 3 + 6)/3 - 2 right ?

OpenStudy (anonymous):

now lets work it out .. (15 * 3 + 6)/3 - 2 (15 * 3 + 6)/3 = 15 + 2 15 + 2 - 2 = 15

OpenStudy (anonymous):

When working it out, for the second step, why does it equal 15 + 2?

OpenStudy (anonymous):

without the subtraction of the 2.. i just looked at this part: (15 * 3 + 6)/3 (15 * 3 + 6)/3 = (15*3)/3 + 6/3 = 15 + 2

OpenStudy (anonymous):

I'm going to eat right now. I can't really focus at the moment. I'll message you if I have any questions.

OpenStudy (anonymous):

ok enjoy

OpenStudy (anonymous):

Thanks, and I understand what you wrote. :)